OpenStudy (anonymous):

14a^2/10b^2 divided by 21a^2/15b^2

OpenStudy (anonymous):

\[\frac{14a^2}{10b^2}*\frac{15b^2}{21a^2}=\frac{210a^2b^2}{210a^2b^2}\]\[=1\]